OpenStudy (anonymous):

Find the length of the rectangle where the length is 12 units more than the width and the perimeter is 8 times the width.

OpenStudy (raden):

perimeter = 8w 2(l+w) = 8w 2(w+12+w) 8w 2w+24+2w=8w solve for w.... then subtitute back w to l=w+12

OpenStudy (anonymous):

so when i substitute w to l=w+12, i don't understand what to do after that...

OpenStudy (anonymous):

Solve the following for w and L, the width and length respectively:\[\{L=12+w,2L+2w=8w\}\]\[\{w=6,L=18\} \]

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

2w+24+2w=8w 4w+24=8w 24 = 8w-4w 24=4w w=24/4=6 l=w+12=6+12=18

OpenStudy (anonymous):

oh okay, thank you(: @RadEn